An inductor has a reactance of 20 Ω and resistance of 25 Ω at a frequency of 60 Hz. What is the inductance of this coil?
1
Expert's answer
2020-12-21T11:02:07-0500

"R=wL\to L=\frac{R}{2\pi\nu}=\frac{20}{2\cdot3.14\cdot60}=0.053(H)"